### v1

# Turbulence to Order Transitions in Activity Patterned Active Nematics

Abstract
We numerically study two-dimensional active nematics with periodic activity patterning. For stripes of activity, we observe a transition from two-dimensional to one-dimensional active turbulence as the maximum active force and distance between activity stripes increases, followed by a transition to stable vortices ordered antiferromagnetically along the stripes and ferromagnetically transverse to the stripes. If the stripes are replaced by columns of activity circles, the vortex ordering is lost. A triangular lattice of activity circles can induce the formation of two-dimensional active turbulence. Our results provide a mechanism for inducing non-equilibrium phase transitions in active nematics using activity inhomogeneity, which can be exploited to create activity patterned ordered phases.
In non-equilibrium active fluids, energy consumed at the microscale generates macroscopic flows [1, 2]. Active nematics, where the constituents and forces are also endowed with apolar anisotropy, appear in cytoskeletal filament suspensions with molecular motors [3, 4, 5, 6], motile bacteria dispersed in passive nematic liquid crystals [7, 8], cellular tissues [9, 10], colonies of elongated soil bacteria [11], and acoustically actuated lyotropic chromonic liquid crystals [12]. In large, unconfined active nematics, the system exhibits “active turbulence,” where the flows are chaotic and topological defect pairs spontaneously unbind [13, 14, 15, 16, 17, 18]. There has been growing interest in developing methods to control the flows and phase behavior of active nematics by changing the substrate friction or external boundaries [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. A particularly promising approach is activity patterning, recently proven achievable using light activated motor proteins dispersed in cytoskeletal filament suspensions [37, 38, 39, 40]. Recent theoretical and numerical studies indicate that activity gradients can separate and orient topological defects and induce multiple active regimes [41, 29, 31, 34].
Here we show that introducing periodic active and inactive stripes or spots can produce transitions from two-dimensional (2D) to one-dimensional (1D) active turbulence or a stable vortical flow state as the activity strength and distance between stripes are varied. Surprisingly, we find that the transitions to turbulence shift to higher active region densities for higher activity. As the active regions become less dense, an ordered vortex state is more likely to appear, with antiferromagnetic ordering along the active channels and ferromagnetic ordering transverse to them. Although vortices still appear if the activity stripes are replaced by columns of activity spots, the ordering of the vortices is lost. Our results provide a new mechanism for stabilizing ordered vortices in active nematics and controlling transitions to and from different active turbulence regimes. Crucially, we find that the activity does not need to be homogeneously distributed in order to produce active turbulence, and that the density of activity itself can control the phase behavior of an active nematic system.
Methods– We model an active nematic using the Beris-Edwards equations for the tensor order parameter $\mathbf{Q}$ [42, 43] and a modified Stokes equation for the incompressible fluid velocity $\mathbf{v}$ [44, 6]:
$\frac{\partial\mathbf{Q}}{\partial t}+\left(\mathbf{v}\cdot\nabla\right)\mathbf{Q}$ $=\mathbf{S}-\frac{\delta F}{\delta\mathbf{Q}}$ (1)
$\ell_{h}^{2}\nabla^{2}\mathbf{v}-\mathbf{v}$ $=\nabla p+(1+\ell_{h}^{2})\nabla\cdot\left(\alpha\mathbf{Q}\right)$ (2)
$\nabla\cdot\mathbf{v}$ $=0.$ (3)
Here $\mathbf{S}$ is a generalized tensor advection, $F$ is a Landau-de Gennes free energy, $\ell_{h}$ is the hydrodynamic screening length characterized by the ratio of fluid viscosity and substrate friction, $p$ is the fluid pressure which is determined from the incompressibility constraint, and $\alpha=\alpha(x,y)$ is the inhomogeneous strength of active forces. All lengths are measured in units of the nematic correlation length and times in units of the nematic relaxation time [45].
To solve Eqs. (1) and (2), we use the finite element method from the C++/Matlab package FELICITY [46]. We consider rectangular domains of various sizes with different patterns of activity $\alpha(x,y)$ set to a maximum (minimum) value of $\alpha=\alpha_{0}$ ($\alpha=0$), and use periodic boundary conditions in all cases. Figure 1(a) shows a system of activity stripes oriented along the $y$ direction, where the active (inactive) regions are of length $\ell_{s}$ ($\ell_{i}$). We fix $\ell_{s}=\ell_{h}=10$ so that the screening length is commensurate with the active stripe length, causing the active stripes to behave as channels of active nematics that interact across an inactive region. We also consider active circles of radius $R=10/\sqrt{\pi}$ arranged in either a triangular lattice [Fig. 1(b)] with varied lattice constant $a$, or vertical lines [Fig. 1(c)] with varied inactive length $\ell_{i}$.
Figure 1: Height maps of activity profiles showing activity level $\alpha$. (a) Active stripes with active region width $\ell_{s}$ and inactive region width $\ell_{i}$. (b) Active circles of radius $R$ arranged in a triangular lattice with lattice constant $a$. (c) Active circles of radius $R$ arranged along vertical lines with inactive width $\ell_{i}$.
Results– We find two transitions in the active stripe array as we vary $\ell_{i}$ and $\alpha_{0}$. For small $\ell_{i}$ and $\alpha_{0}$, the behavior qualitatively resembles the active turbulence that would appear if the activity were homogeneous [4, 47, 6, 48], as shown by the plot of the time-averaged normalized vorticity $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ and velocity $\langle\mathbf{v}\rangle$ fields for a system with $\ell_{i}=6$ and $\alpha_{0}=0.7$ in Fig. 2(a). Both quantities have a value close to 00 throughout the domain, and there is no correlation between the active regions and the weak vortical patterns or the time-averaged topological defect density [49, 50, 45]. When either $\alpha_{0}$ or $\ell_{i}$ is increased, as shown in Fig. 2(b) at $\ell_{i}=16$ and $\alpha_{0}=0.7$ and in Fig. 2(b) at $\ell_{i}=6$ and $\alpha_{0}=1.2$, a charge separating behavior arises. Here $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ becomes strongly nonzero only in the active regions while $|\langle\mathbf{v}\rangle|$ is suppressed in the active regions, so that steady flows occur only in the inactive regions between the active stripes. The time-averaged topological defect density separates, and positive topological defects occupy the active regions while negative topological defects remain in the inactive regions [45]. A similar charge separation was recently predicted theoretically and observed numerically for a single stripe of activity [34]. At large $\alpha_{0}$, the system develops stable and sometimes static vortices, as illustrated in Fig. 2(d) at $\ell_{i}=16$ and $\alpha_{0}=1.5$, where $\langle\mathbf{v}\rangle$ has clear vortical signatures. The nematic configuration in this regime consists of two positive defects rotating around one another near the center of a given vortex, while negative defects sit in between the vortices [45].
Figure 2: (a-d) Spatial plots of scaled time-averaged vorticity field $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ (color) with time-averaged velocity field $\langle\mathbf{v}\rangle$ (arrows), where bright (dark) areas indicate $\alpha=\alpha_{0}$ ($\alpha=0$). (a) $\ell_{i}=6$ and $\alpha_{0}=0.7$. (b) $\ell_{i}=16$ and $\alpha_{0}=0.7$. (c) $\ell_{i}=6$ and $\alpha_{0}=1.2$. (d) $\ell_{i}=16$ and $\alpha_{0}=1.5$. Animations of the dynamics of each state are available in the Supplemental Material [45]. (e) Heat map of the logarithm of the excess kurtosis of the vorticity distribution, $\log|K|$, over the whole domain as a function of $\alpha_{0}$ vs $\ell_{i}$. (f) Heat map of $\log|K|$ only measured over active regions as a function of $\alpha_{0}$ vs $\ell_{i}$. Dashed line: points from panel (e) where $\log|K|=0$. Solid line: points from panel (f) where $\log|K|=0$. Letters: locations at which images in panels (a-d) were obtained.
To quantify the transitions, we compute the excess kurtosis of the vorticity distributions $P(\omega)$:
$K[\mathcal{P}(\omega)]\equiv\frac{\mu_{4}}{\sigma^{4}}-3$ (4)
where $\mu_{4}$ is the fourth moment of $P(\omega)$ and $\sigma$ is the standard deviation of $P(\omega)$. As a statistical tool, $K$ characterizes the size of tails of a distribution, with $K=0$ for a Gaussian distribution and $K<0$ ($K>0$) for distributions with smaller (larger) tails. We use $K$ as a measure for active turbulence since it has been shown both computationally and experimentally that active turbulence in active nematics is associated with Gaussian vorticity distributions [51, 48, 31]. In Fig. 2(e) we plot a heat map of $\log|K|$ measured over the entire spatial domain as a function of $\alpha_{0}$ versus $\ell_{i}$. The dashed line indicates points where $\log|K|=0$, and we define the $\log|K|<0$ region below this line to be in 2D active turbulence. Next, we perform a measurement of $K[\mathcal{P}(\omega|\alpha>0.5\alpha_{0})]$, in which we restrict the domain to only the active portions of the sample and exclude the inactive areas. We plot the resulting heat map of $\log|K|$ as a function of $\alpha_{0}$ versus $\ell_{i}$ in Fig. 2(f). The solid line indicates points for which $\log|K|=0$. In the window between the solid and dashed lines, where the active regions produce a turbulent signature but the inactive regions do not, 1D active turbulence has emerged along each active stripe, but turbulent behavior is suppressed in the regions between the stripes. Thus, the two transitions in the system are between 2D and 1D active turbulence, and between 1D turbulence and a vortex state.
Figure 3: (a,b) Time snapshots of $\omega/|\omega|_{max}$ (color) and ${\bf v}$ (arrows) for a triangular lattice of active circles (bright spots). (a) $a/2R=1.28$ and $\alpha_{0}=0.75$. (b) $a/2R=1.74$ and $\alpha_{0}=1.25$. Animations of the dynamics of each state are available in the Supplemental Material [45]. (c) Heat map of $\log|K|$ as a function of $\alpha_{0}$ vs $a/2R$. Line: points where $\log|K|=0$. Letters: locations at which images in panels (a,b) were obtained.
To further probe the turbulent to vortex transition, we simulate a system in which the activity is confined to circles that are arranged in a triangular lattice, as shown in Fig. 1(b). The area fraction of the active regions is given by $\nu=(\pi/2\sqrt{3})(2R/a)^{2}$, where $R$ is the circle radius and $a$ is the lattice constant. We vary $a/2R$ to give values of $\nu$ in the range $\nu=0.1$ to $\nu=0.7$. Even without the 1D stripe arrangement of activity, we observe a transition from active turbulence to a stable vortex state when $\alpha_{0}$ and $a/2R$ are increased. In Fig. 3(a) we plot a time snapshot of the vorticity and velocity fields for a simulation with $a/2R=1.28$ and $\alpha_{0}=0.75$. Here we observe active turbulence in which the locations of the vortices are not correlated with the locations of the activity spots. When we increase $\alpha_{0}$ and $a/2R$, as shown in Fig. 3(b) for $\alpha_{0}=1.25$ and $a/2R=1.74$, individual vortices become localized in each active region. In Fig. 3(c), we plot a heat map of $\log|K|$ measured over the entire sample as a function of $\alpha_{0}$ versus $a/2R$. Below the line marking points at which $\log|K|=0$ the system exhibits active turbulence, and above the line the system relaxes to a configuration with a stable vortex at each active site. The turbulent to vortex transition has a similar form to what we observed in the active stripe sample from Fig. 2(e,f), but the 1D active turbulent state is lost due to the geometry of the active circles.
It is surprising that for both the active stripes and the active circles, the transition to the turbulent state shifts to smaller spacings between the active regions (and a greater total active area) when the activity level $\alpha_{0}$ is increased since, in confined systems with homogeneous activity, an increase in the activity typically produces an increased amount of topological defect unbinding and a more rapid transition to turbulence [4, 51, 6]. In our systems with inhomogeneous activity, we can define an active length scale $\ell_{\alpha}\propto\sqrt{1/\alpha_{0}}$ and a characteristic length scale $\ell_{s}$ of the active regions. We find that when $\ell_{\alpha}\gtrsim\ell_{s}$, the system exhibits active turbulence, while when $\ell_{\alpha}\lesssim\ell_{s}$, localized vortical structures appear [45]. This has implications even for systems that do not have explicitly patterned activity, since in any experimental system the activity is not distributed homogeneously. Our results indicate that the approximation of homogeneously distributed activity is valid in the limit of small $\ell_{s}$ and small $\alpha_{0}$, but outside this limit, the effect of inhomogeneous activity may disrupt the active turbulent phase.
A more careful inspection of the vortex state that appears in the active stripe system for large $\alpha_{0}$, as in Fig. 2(d), reveals antiferromagnetic ordering of the vortices along the $y$ direction and ferromagnetic ordering along the $x$ direction. To quantify this ordering, we compute the autocorrelation functions of the vorticity along the $x$ and $y$ directions at various values of $y_{0}$ and $x_{0}$:
$C_{x}(X,y_{0},t)$ $=\frac{\langle\omega(x+X,y_{0},t)\omega(x,y_{0},t)\rangle_{x}}{\langle\omega(x,y_{0},t)^{2}\rangle_{x}}$ (5)
$C_{y}(x_{0},Y,t)$ $=\frac{\langle\omega(x_{0},y+Y,t)\omega(x_{0},y,t)\rangle_{y}}{\langle\omega(x_{0},y,t)^{2}\rangle_{y}}.$ (6)
We define $\chi_{x}(t)=\langle C_{x}(\ell_{s}+\ell_{i},y_{0},t)\rangle_{y_{0}}$ and $\chi_{y}(t)=\langle C_{y}(x_{0},\ell_{h},t)\rangle_{x_{0}}$ so that $\chi_{x}$ ($\chi_{y}$) measures the nearest neighbor correlation in the $x$ ($y$) direction. Here, $\chi_{x,y}\in[-1,1]$, with positive (negative) values indicating ferromagnetic (antiferromagnetic) order and larger magnitudes indicating stronger order. An example of the calculation of $\chi_{x}$ and $\chi_{y}$ appears in the Supplemental Material [45].
Figure 4: (a, b) Vortex correlation parameters $\chi_{x}$ and $\chi_{y}$ versus time. (a) $\ell_{i}=20$ and $\alpha_{0}=1.5$. (b) $\ell_{i}=8$ and $\alpha_{0}=1.4$. (c) Heat map of the time-averaged $\langle\chi_{x}\rangle$ as a function of $\alpha_{0}$ vs $\ell_{i}$. (d) Heat map of the time-averaged $\langle\chi_{y}\rangle$ as a function of $\alpha_{0}$ vs $\ell_{i}$. Letters: locations at which curves in panels (a, b) were obtained.
We compute $\chi_{x,y}$ for $\alpha_{0}>1.2$, where $\log|K|$ is large and positive. In Fig. 4(a) at $\ell_{i}=20$ and $\alpha_{0}=1.5$, for long times $\chi_{x}\approx+0.5$ and $\chi_{y}\approx-1$, indicating moderate ferromagnetic order along the $x$ direction and strong antiferromagnetic order along the $y$ direction. For smaller $\ell_{i}$, $\chi_{x}$ and $\chi_{y}$ fluctuate more over time, as shown in Fig. 4(b) at $\ell_{i}=8$ and $\alpha_{0}=1.4$. Here, $\chi_{x}$ fluctuates around zero but $\chi_{y}$ fluctuates around $\chi_{y}=-0.5$, indicating a lack of ordering along the $x$ direction and moderate antiferromagnetic order along the $y$ direction. In Figs. 4(c,d) we plot heat maps of the time averaged $\langle\chi_{x}\rangle$ and $\langle\chi_{y}\rangle$, respectively, as a function of $\alpha_{0}$ versus $\ell_{i}$. As $\ell_{i}$ increases, $\langle\chi_{x}\rangle$ grows from roughly zero to $\approx 0.6$, indicating the emergence of ferromagnetic order transverse to the activity stripes. In contrast, $\langle\chi_{y}\rangle$ is almost always negative, and reaches values very close to $-1$ when $\ell_{i}>14$, indicating almost perfect, stable antiferromagnetic order along the active stripe.
Antiferromagnetic vortex ordering has been observed for active nematics confined to channels, as well as in two-dimensional systems with increasing substrate friction or obstacle confinement [51, 52, 22, 25, 32, 36]. The unbinding of topological defects helps promote antiferromagnetic order when vortices are stable, since the negative defects can sit between the vortices and mediate a rotation of the nematic director as the chirality of the positive defects at the center of each vortex alternates. A time average of the topological charge density in our active stripe system reveals negative defects sitting between the vortices for simulations in the large $\ell_{i}$ and large $\alpha_{0}$ regime [45]. The ferromagnetic order transverse to the active regions is of great interest since, to our knowledge, it has only been reported numerically for active nematics interacting with periodic arrays of obstacles [32], while vortex synchronization has only recently been reported in acoustically actuated active nematics [12]. For well separated active stripes, the only interaction that couples the stripes is the elastic interaction of the nematic, which has been shown to promote ferromagnetic order of vortices in obstacle laden systems [32]. We therefore conclude that the ferromagnetic order transverse to the stripes is driven by elastic interactions, while the antiferromagnetic order along the stripes is driven by active forces.
Interestingly, we find that separation of active regions is necessary for strong vortex ordering. When the stripes are closer together, flow generation between the stripes disrupts the vortex order, as illustrated in the Supplemental Movie 7 [45], leading to the fluctuations seen in Fig. 4(b) and bringing $\langle\chi_{x,y}\rangle$ closer to zero. These results are more in line with with earlier numerical results on activity patterns with no gaps in the activity [31].
Figure 5: (a) Time snapshot of $\omega/|\omega_{max}|$ (color) and $\langle{\bf v}\rangle$ (arrows) for lines of activity circles (bright spots) with $\ell_{i}=14$ and $\alpha_{0}=1.5$. (b) The corresponding time-averaged topological defect density. (c) $\langle\chi_{x}\rangle$ and (d) $\langle\chi_{y}\rangle$ vs $\ell_{i}$ for systems of activity stripes (blue) and circle lines (red) with $\alpha_{0}=1.5$.
Finally, we investigate the impact of the 1D degree of freedom along the stripes on the vortex ordering. While the antiferromagnetic order develops quickly along the active stripes, the ferromagnetic order takes much longer to develop. Qualitatively, the vortices slide into place over time in order to develop ferromagnetic order along the $x$ direction, rather than nucleating in ordered positions initially. To remove the 1D degree of freedom, we simulate lines of active circles [Fig. 1(c)] for $\alpha_{0}=1.5$ at varied $\ell_{i}$. As shown in Fig. 5(a), vortices quickly develop in the active regions, while Fig. 5(b) indicates that positive topological defects rotate in the center of the vortices while negative defects sit between active regions. Although this is similar to the behavior of the active stripe system, the vortex ordering is different. In Figs. 5(c) and (d) we plot $\langle\chi_{x}\rangle$ and $\langle\chi_{y}\rangle$ versus $\ell_{i}$ for the stripe and circle line systems. The measures diverge for large $\ell_{i}$, with $\langle\chi_{x}\rangle\approx 0$ and $\langle\chi_{y}\rangle\gtrsim 0$ for the circle lines, but both measures reaching nonzero values for the stripe system. The vortices develop quickly due to the large active forces in the active regions, but in the circle line system, they remain frozen in place and cannot reorder along $x$ [45]. Ordinarily the vortices would naturally adopt an antiferromagnetic ordering along $y$ in a 1D channel, as in the stripe system, but this ordering requires an odd number of negative topological defects to sit between pairs of vortices to mediate the change in chirality. As shown in Fig. 5(b), in the circle line system, the divots between adjacent circles permit two defects to sit between vortices along the edge of the inactive region. We find that when one defect is present, the chirality of neighboring vortices changes, but when two defects are present, the chirality remains the same. This result underscores the pivotal role that activity geometry plays in determining the ordering of the system.
Conclusion– We have studied the effects of periodic, inhomogeneous activity patterning on active nematics. We observe two non-equilibrium transitions for a system of activity stripes: a transition from 2D to 1D active turbulence, and a transition from 1D active turbulence to a well-defined vortical flow state that can exhibit ferromagnetic and antiferromagnetic vortex ordering. The vortex ordering can be destroyed if the activity stripes are replaced by lines of activity circles. When we introduce a triangular array of activity circles, the transition to a turbulent state is suppressed if the activity is too large due to the tendency of localized vortices to emerge at the activity spots. Our results provide a method for inducing and controlling non-equilibrium phase transitions in active nematics and for producing customized activity patterned ordered states.
Acknowledgements.
We gratefully acknowledge the support of the U.S. Department of Energy through the LANL/LDRD program for this work. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 89233218CNA000001). CDS also gratefully acknowledges support through the William H. Miller III fellowship from the Department of Physics and Astronomy at Johns Hopkins University.